One type of cat food recommends that a cat have a daily serving of 1/3ounce of dry cat food per pound of body weight. If a cat is fed 3 2/3 ounces of dry food a day according to these recommendations, how many pounds does the cat weigh?
Found 2 solutions by Nate, Shin123:
Answer by Nate(3500)

x(1/3) = 3+(2/3)
x(1/3) = 11/3
x = 11
The cat weighs a whopping eleven pounds.

Answer by Shin123(626)